OpenStudy (anonymous):

Find the radius of a cone with volume 199.23 cm3 and height 9 cm.

OpenStudy (anonymous):

@mathmale mind helping with another :)

OpenStudy (nikato):

U can look here to find formula for volume of cone http://math.about.com/od/formulas/ss/surfaceareavol_2.htm

OpenStudy (nikato):

Plug in 199.23 for V And 9 for h. And then try to solve for r

OpenStudy (anonymous):

V=1/3 * pie * r square * h Plug in the variables you have and sole for r Don't forget to square root it

OpenStudy (anonymous):

ok im lost as to how to eliminate pi...

OpenStudy (nikato):

So 199.23=1/3(pi)(r^2)(9) 199.23 =3(pi)(r^2) -------- -------- 3pi. 3pi 66.41 ------ = r^2 pi Then square root both sides to find r

OpenStudy (nikato):

It'll be better to do it with a calculator

OpenStudy (anonymous):

so its 4.5 :)

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

Yea. Around there 4.5 or 4.6